The Bigbee Bottling Company is contemplating the replacement of one of its bottling machines with a newer and more efficient one. The old machine was purchased prior to the new tax legislation, has a book value of $600,000, and a remaining useful life of 5 years. The firm does not expect to realize any return from scrapping the old machine in 5 years, but it can sell it now to another firm in the industry for $265,000. The old machine is being depreciated by $120,000 per year, using the straight-line method.

The new machine has a purchase price of $1,175,000, an estimated useful life of 5 years, and an estimated salvage value of $145,000. The new machine is eligible for 100% bonus depreciation at the time of purchase. It is expected to economize on electric power usage, labor, and repair costs, as well as to reduce the number of defective bottles. In total, an annual savings before taxes of $220,000 will be realized if the new machine is installed. The company’s marginal tax rate is 25% and it has a 12% WACC.

  1. What initial cash outlay is required for the new machine after bonus depreciation is considered?

  2. Calculate the change in the annual depreciation expense if the replacement is made.

  3. What are the incremental cash flows in Years 1 through 5?

  4. Should the firm purchase the new machine? Support your answer.

  5. In general, how would each of the following factors affect the investment decision, and how should each be treated?

    1. The expected life of the existing machine decreases.

    2. The WACC is not constant, but is increasing as Bigbee adds more projects into its capital budget for the year.
Answer #1
Initial Cash Outlay
Purchase price of New Machine A 1175000
100% Bonus depreciation B 1175000
Tax Rate C 25%
Tax Benefit D=BXC 293750
Sales value of old machine E 265000
Net cash Outlay F=A-D-E 616250
As sales value of old machine is less than book value hence no tax implication arise on the sales realization.
Change in Per annum Depreciation
a) If Firm avail 100% Bonus Depreciation option.
Year Depreciation
Old Machine New Machine Change
A B C D=B-C
1 120000 1175000 -1055000
2 120000 120000
3 120000 120000
4 120000 120000
5 120000 120000
Total 600000 1175000 575000
a) If Firm does not avail 100% Bonus Depreciation option.
Year Depreciation
Old Machine New Machine Change
A B C D=B-C
1 1175000 206000 969000
2 1175000 206000 969000
3 1175000 206000 969000
4 1175000 206000 969000
5 1175000 206000 969000
Total 5875000 1030000 1030000
Calculation of incremental Cash Flow
Cash Flow in case of new Machine
Year 1 2 3 4 5 Total
Purchase Cost A -1175000 -1175000
Sale of Old Machine B 265000 265000
Saving in Expenses C= Given 220000 220000 220000 220000 220000 1100000
Change in Depreciation D =calculated Above -1055000 120000 120000 120000 120000 -575000
Expenses allowable for tax Deduction E=c+d -835000 340000 340000 340000 340000 525000
Tax Expenses F=E X 25% -208750 85000 85000 85000 85000 131250
Net incremental expenses G=E-F -626250 255000 255000 255000 255000 393750
Net Cash Flow H=A+B-D+G -481250 135000 135000 135000 135000 58750
Cost of Capital I 1.0000 0.8929 0.7972 0.7118 0.6355 0.5674
Net Present Value J=HXI -4,81,250.00 1,20,535.71 1,07,621.17 96,090.33 85,794.94 33,336.33
NPV is positive hence should purchase new machine.
Different conditions
1) When expected life of existing machine is decreasing
As old machine's life is decreasing it is better to replace with new machine as early as possible
because realizable value of old machine can be decrease further which can reduce benefits to the company.
2)when WACC is not constant and increasing
Should not replace its old machine as margin is very low and high cost of capital can give negative results.